# Advection of the image point in probabilistically-reconstructed phase spaces

### v1

@@ 3.3 NEMO model
The dimensionality of the probabilistically-reconstructed space is $n\approx 30000$. One can run the HP method in the high-dimensional phase space, but in this work we further reduce it down to $m=200$ by using the EOF-PC decomposition (Preisendorfer, 1988). It results in a two-order of magnitude reduction compared with the original dimension; note that the 200 leading EOFs capture 99% of the total SST variance and 60% of the total SRV variance. We compute the EOF-PC decomposition of SST and SRV separately. However, it can be computed for both if one needs to model these fields together.
The space of PCs (let it be denoted as $\mathbf{x}$) can be regarded as a reduced version of the probabilistically-reconstructed phase space. Given this reduced space, one can reconstruct a dynamical system governing the reduced dynamics of observed reference data, as we did in (Shevchenko and Berloff, 2022b). In this study, we use a different idea: the dynamics of PCs is modelled in the reduced space with the HP method directly and then the HP solution is projected back into the full-dimensional probabilistically-reconstructed phase space.
@@
$\frac{d\mathbf{z}}{dt}=\frac{1}{M}\sum\limits_{j=\mathcal{U}_{J}}\mathbf{F}(\mathbf{x}_{j})\Big{|}_{\mathcal{U}(\mathbf{z}(t))}+\eta\left(\frac{1}{N}\sum\limits_{i=\mathcal{U}_{I}}\mathbf{x}_{i}\Big{|}_{\mathcal{U}(\mathbf{z}(t))}-\mathbf{z}(t)\right),\quad\mathbf{z}(t_{0})=\mathbf{z}_{0}.$ (4)
The only difference of (4) compared with equation (1) is that $\mathbf{x}$ does not represent the reference solution (it would be SST or SRV in this case) but the PCs. Having solved equation (4), we approximate the HP solution, $\mathbf{y}(t)$, in the probabilistically-reconstructed phase space by using the leading EOF-PC pairs as follows:
@@
We report the results in Fig. 10 which clearly shows that the HP solution (Fig. 10b) much better represents the Gulf Stream than the modelled solution (Fig. 10c). These results are also confirmed in Fig. 11. Namely, the difference between the time mean of the reference solution and HP solution is much smaller (Fig. 11a) than that of the modelled solution (Fig. 11b). The root mean square error for the HP solution (Fig. 11c) is also much lower then the one of the modelled solution (Fig. 11d).
@@
Figure 11: Shown are (a) difference between the time mean reference SST and HP solution, (b) difference between the time mean reference and modelled SST, (c) root mean square error between the reference SST and HP solution, (d) root mean square error between the reference and modelled SST. We use $m=200$ in the EOF-PC decomposition.
More insights about how the HP solution differs from the modelled solution can be seen in the SRV fields (Fig. 12). The Gulf Stream computed with the HP method (Fig. 12b) is much stronger than the one computed with the 1/4 model (Fig. 12c). Moreover, the HP solution is teemed with vortices (like the reference one), while the vortex dynamics is significantly inhibited in the modelled solution as it is also reflected by the error plots in Fig. 13. Doubling the number of PCs ($m=400$) results in a more pronounced Gulf Stream dynamics and larger population of vortices (Fig. 12). However it is contribution to the time mean flow and the root mean square error is small (not shown).
@@
Figure 13: The same as in Fig. 11 but for SRV.
@@ 4 Conclusions
In this study we have proposed to use the joint probability distribution (JPD) to probabilistically reconstruct the phase space which can be used when the original reference phase space (computed from numerical solutions, observations, or both) is too sparse to allow the use of data-driven methods. The probabilistically-reconstructed phase space is calculated by sampling from the JPD, thus providing an extra source of data from the reference distribution and ensuring that both the reference phase space and the reconstructed one share the same reference distribution. We have shown how the probabilistic reconstruction works on the example of Chua’s system and then applied it within the context of the global ocean model NEMO. We have also found that the HP method can be used directly in reduced phase spaces thus allowing several-orders-of-magnitude acceleration compared to the 1/4-resolution NEMO run. We reduced the dimensionality of the reference phase space by two orders of magnitude by applying the EOF-PC decomposition.
Our results show that the use of the HP method in the probabilistically-reconstructed (and reduced) phase space gives more accurate results and more realistic flow dynamics compared with the 1/4-NEMO simulation. It is also worth mentioning that the two order of magnitude acceleration can be translated into using larger ensemble of HP solutions for probabilistic predictions.
We believe that the proposed method demonstrates the strong potential for the use in operational ocean and ocean-atmospheric models. The next step in this direction is to assess prediction skills and achievable forecast range of the proposed method for high-resolution forecasting systems and near-term climate predictions.
Acknowledgements.
[…]

### v2

@@ 3.2 Calibration of hyper-parameters
3.3 Corrupted reference data
Corrupted reference data is ubiquitous in the realm of data-driven methods. Reference data sets lacking states in the phase space or having vast voids (regions of no data) can significantly influence the accuracy of data-driven methods or make them inapplicable. Therefore, it is instructive to study what happens when the reference data is damaged. As an example, we consider three different cases of the reference phase space corrupted with: (1) randomly-missed states, (2) gaps, and (3) cuts (Fig. 9). In all cases, sampling from the JPD reconstructs the phase space up to the extent, which is enough for the HP method to work. In the first case, sampling from the JPD densifies the attractor well (top subplot in Fig. 9b). In the second and third cases, however, there are still some voids in the reconstructed phase space (middle and bottom subplots in Fig. 9b) which cannot be completely filled in by sampling from the JPD, but the HP method does reproduce the dynamics on the attractor (similar to the reference one) nonetheless.
Figure 9: Shown are (a) reference data sets with different level of damage (top – (1) 50% of randomly-chosen states has been removed from the original reference data shown in Fig. 8a, middle – (2) original reference data with gaps, bottom – (3) original reference data with three cuts); (b) phase spaces probabilistically-reconstructed from the corresponding corrupted reference data sets; (c) HP solutions computed in the probabilistically-reconstructed phase spaces. In all cases, the HP method can reproduce the solution on the attractor even when sampling from the JPD cannot completely repair the damage. Note that for the HP method we use $\eta=0.003$ for cases (1) and (2), and $\eta=0.051$ for case (3).
The ability of the HP method to work in corrupted phase spaces can be exploited as a fast reanalysis method in which the complex dynamics of a comprehensive ocean model is replaced with the HP solution. The method can be thought of as a dynamic interpolation which can stitch gaps in observational data with the HP dynamics.
@@ 3.4 NEMO model
The dimensionality of the probabilistically-reconstructed space is $n\approx 30000$. One can run the HP method in the high-dimensional phase space, but in this work we further reduce it down to $m=200$ by using the EOF-PC decomposition [28]. It results in a two-order of magnitude reduction compared with the original dimension; note that the 200 leading EOFs capture 99% of the total SST variance and 60% of the total SRV variance. We compute the EOF-PC decomposition of SST and SRV separately. However, it can be computed for both if one needs to model these fields together.
The space of PCs (let it be denoted as $\mathbf{x}$) can be regarded as a reduced version of the probabilistically-reconstructed phase space. Given this reduced space, one can reconstruct a dynamical system governing the reduced dynamics of observed reference data, as we did in [23]. In this study, we use a different idea: the dynamics of PCs is modelled in the reduced space with the HP method directly and then the HP solution is projected back into the full-dimensional probabilistically-reconstructed phase space.
@@
$\frac{d\mathbf{z}}{dt}=\frac{1}{M}\sum\limits_{j{\color[rgb]{0,0,0}\in}\mathcal{U}_{J}}\mathbf{F}(\mathbf{x}_{j})+\eta\left(\frac{1}{N}\sum\limits_{i{\color[rgb]{0,0,0}\in}\mathcal{U}_{I}}\mathbf{x}_{i}-\mathbf{z}(t)\right),\quad\mathbf{z}(t_{0})=\mathbf{z}_{0},$ (4)
where $\mathcal{U}_{I}$ and $\mathcal{U}_{J}$ are the sets of timesteps indexing the discrete reference solution $\mathbf{x}$ in the neighbourhood of $\mathbf{z}(t)$. The only difference of (4) compared with equation (1) is that $\mathbf{x}$ does not represent the reference solution (it would be SST or SRV in this case) but the PCs. Having solved equation (4), we approximate the HP solution, $\mathbf{y}(t)$, in the probabilistically-reconstructed phase space by using the leading EOF-PC pairs as follows:
@@
We report the results in Fig. 11 which clearly shows that the HP solution (Fig. 11b) much better represents the Gulf Stream than the modelled solution (Fig. 11c). These results are also confirmed in Fig. 12. Namely, the difference between the time mean of the reference solution and HP solution is much smaller (Fig. 12a) than that of the modelled solution (Fig. 12b). The root mean square error for the HP solution (Fig. 12c) is also much lower then the one of the modelled solution (Fig. 12d). The time mean energy spectral density of the HP solution is closer to the reference than that of the modelled solution over the whole range of wave numbers (Fig. 12e, compare the blue and magenta graphs with the black one), and this difference is even more pronounced for SRV (Fig. 14e).
@@
Figure 12: Shown are (a) difference between the time mean reference SST and HP solution, (b) difference between the time mean reference and modelled SST, (c) root mean square error between the reference SST and HP solution, (d) root mean square error between the reference and modelled SST, (e) the natural logarithm of the time mean energy spectral density $\langle ESD\rangle$ vs the magnitude of the wave vector $\mathbf{k}$. We use $m=200$ in the EOF-PC decomposition for (a)-(d). The $\langle ESD\rangle$ of the HP solution (red and magenta) is closer to the reference (black) than that of the modelled solution (blue). Note that $\langle ESD\rangle$ for $m=200$ (red) and $m=400$ (blue) are almost the same, as adding more EOFs does not contribute much to the accuracy of SST. For illustrative purposes, we apply a locally estimated scatterplot smoothing (also known as a Savitzky–Golay filter) to smooth the ESD [29].
More insights about how the HP solution differs from the modelled solution can be seen in the SRV fields (Fig. 13). The Gulf Stream computed with the HP method (Fig. 13b) is much stronger than the one computed with the 1/4 model (Fig. 13c). Moreover, the HP solution is teemed with vortices (like the reference one), while the vortex dynamics is significantly inhibited in the modelled solution as it is also reflected by the error plots in Fig. 14, as well as by the time mean energy spectral density, denoted as $\langle ESD\rangle$. The latter is more sensitive to the number of EOFs compared with SST, i.e. the more EOFs are used, the more accurate the $\langle ESD\rangle$ is. It is because SRV fields contain much more small-scale features (compared with SST), which therefore require more EOFs to be accurately represented. Doubling the number of PCs ($m=400$) results in a more pronounced Gulf Stream dynamics and larger population of vortices (Fig. 13), as well as in a more accurate representation of $\langle ESD\rangle$ (Fig. 14e). However, it is contribution to the time mean flow and the root mean square error is small (not shown).
@@
Figure 14: The same as in Fig. 12 but for SRV. Note that adding more EOFs makes the time mean energy spectral density, $\langle ESD\rangle$, closer to the reference one; compare the $\langle ESD\rangle$ of the HP solution for $m=200$ (red) and $m=400$ (magenta) with the reference (black). The $\langle ESD\rangle$ of the modelled solution (blue) is much lower compared with that of the HP solution, especially for smaller scales.
@@ 4 Conclusions
In this study we have proposed to use the joint probability distribution (JPD) to probabilistically reconstruct the phase space which can be used when the original reference phase space (computed from numerical solutions, observations, or both) is too sparse to allow the use of data-driven methods. The probabilistically-reconstructed phase space is calculated by sampling from the JPD, thus providing an extra source of data from the reference distribution and ensuring that both the reference phase space and the reconstructed one share the same reference distribution. We have shown how the probabilistic reconstruction works on the example of Chua’s system and then successfully applied it within the context of the global ocean model NEMO.
The current methodology is limited to the phase space of HP solution computed in the probabilistically-reconstructed phase space. The phase space of HP solution is larger than the phase space computed from the data distribution (i.e. the probabilistically-reconstructed one), because the latter might have voids (which cannot be patched up by sampling from the JPD), while the former has no voids. Giving these limitations, the proposed methodology does allow us to compute the HP solution in the probabilistically-reconstructed (and reduced) phase space which is more accurate compared with the 1/4-NEMO simulation nonetheless. However, if the reference distribution changes over time (e.g. the reference solution drifts away from the phase space, used to compute the reference distribution) then the HP solution cannot follow these changes.
We have also found that the HP method can be used directly in reduced phase spaces thus allowing several-orders-of-magnitude acceleration (compared to the 1/4-resolution NEMO run), which can be translated into using larger ensemble of HP solutions for probabilistic predictions. Note that we compare only the degrees of freedom used in the HP method and those used in the NEMO model to get the solution at $1/4^{\circ}$ resolution. It goes without saying that the NEMO model evolves much more variables than the HP method, but in this comparison we pretend that NEMO needs only $\sim$30000 grid points to compute the solution, while the HP method uses only 200 in the compressed phase space, and more importantly gives a solution which is closer to the reference one than that of the $1/4^{\circ}$ NEMO simulation.
We have studied how the proposed method performs on corrupted reference data (for which we considered three cases: randomly-missed states, gaps, and cuts) and found that the method works well even when sampling from the JPD cannot patch up voids in the reference phase space. It allows us to conclude that the HP method in probabilistically-reconstructed phase spaces can be used as a fast reanalysis in which the complex dynamics of a comprehensive ocean model is replaced with the HP solution, as well as a dynamic interpolation method to stitch gaps in observational data.
The primary utility of the proposed methodology is to serve as a fast ocean, atmospheric, or ocean-atmospheric emulator in the reference phase space, or as an alternative to parameterisations in comprehensive ocean and ocean-atmospheric models. We believe that it demonstrates the strong potential and it can be of use in operational ocean and ocean-atmospheric models.
For studying climate change, one might think that the reference data set might be a limiting factor. And indeed, it is, but only if future climate states are not within the reference phase space. As it is a priori unknown whether future climate states are located in the reference phase space, one might consider using this methodology as an alternative to parameterisations in comprehensive models for near-term climate predictions (NTCP), instead of using it as a tool for forecasting, since more research is needed to address this question. Therefore, the next step in this direction would be to assess prediction skills and achievable forecast range of the method alone and in combination with high-resolution forecasting systems for NTCP.
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Data availability
The output of the global 1/4 (ORCA025-N4001) and 1/12 (ORCA0083-N006) resolution NEMO model used in this study is available on JASMIN ( https://jasmin.ac.uk/users/access/ ) from these locations:
/gws/nopw/j04/nemo_vol1/ORCA025-N401 and
/gws/nopw/j04/nemo_vol1/ORCA0083-N006 .
Acknowledgments
The author thanks the Natural Environment Research Council for the support of this work through the projects CLASS and ATLANTIS (P11742-01), as well as Andrew Coward and Chris Wilson for the production of and help with the NEMO datasets, respectively.